Assume that it is currently January 15, 2016. The bonds have a par value of $2,000.

Company - the Corp.

Coupon - 7.200 in percent

Maturity - Jan 15, 2027

Last Price - ??

Last Yield - 5.30 in percent

EST $ Vol (000’s) - 48,949

Current Yield is 6.23%

What price in USD would you expect to pay for the Corp. bond?
Answer #1
                  K = Nx2
Bond Price =∑ [(Semi Annual Coupon)/(1 + YTM/2)^k]     +   Par value/(1 + YTM/2)^Nx2
                   k=1
                  K =11x2
Bond Price =∑ [(7.2*2000/200)/(1 + 5.3/200)^k]     +   2000/(1 + 5.3/200)^11x2
                   k=1
Bond Price = 2313.7
Using Calculator: press buttons "2ND"+"FV" then assign
PMT = Par value * coupon %/coupons per year=2000*7.2/(2*100)
I/Y =5.3/2
N =11*2
FV =2000
CPT PV
Using Excel
=PV(rate,nper,pmt,FV,type)
=PV(5.3/(2*100),2*11,-7.2*2000/(2*100),-2000,)